Find the derivative of the vector function.
a)r(t) =<e^(-t),t-t^3,ln t>
b)r(t) =(1/(1 +t)i+(t/(1 +t))j+((t^2)/(1 +t))k
1
Expert's answer
2020-09-17T15:28:05-0400

(a)

Let's use the [ ] instead of the < > such that e-t

"𝐫 (t)=\begin{bmatrix}
 e^{-t} \\
t-t^3\\
ln(t)

 
\end{bmatrix}" so that the derivative of r (t) will be;


"𝐫'(t)=\frac{d𝐫}{dt} (t) = \begin{bmatrix}
 \frac{d}{dt} e^{-t} \\
\frac{d}{dt} t-t^3 \\
\frac{d}{dt}ln (t)


 
\end{bmatrix}"

"=\begin{bmatrix}
 -e^{-t} \\
1- 3t^3 \\
\frac{1}{t}

 
\end{bmatrix}"


(b)


"𝐫(t)=\frac{1}{1+t}i + \frac{t}{1+t}j + \frac{t^2}{1+t}k" here, we apply the formula;


"𝐫'(t)= \lim\limits_{Δt→0} \frac{𝐫(t + Δt) - 𝐫(t)}{Δt}"


"=\lim\limits_{Δt→0} \frac{\frac{1}{1+t + Δt}i + \frac{t + Δt}{1+t + Δt}j + \frac{(t + Δt)^2}{1+t + Δt}k- (\frac{1}{1+t}i + \frac{t}{1+t}j + \frac{t^2}{1+t}k)} {Δt}"


"=\lim\limits_{Δt→0} \frac{\frac{1}{1+t + Δt}i + \frac{t + Δt}{1+t + Δt}j + \frac{(t^2 + 2tΔt +(Δt)^2}{1+t + Δt}k- (\frac{1}{1+t}i + \frac{t}{1+t}j + \frac{t^2}{1+t}k)} {Δt}"


"=\lim\limits_{Δt→0}\frac{[\frac{1}{1+t + Δt}- \frac{1}{1+t}] i + [\frac{t + Δt}{1+t + Δt}-\frac{t}{1+t}]j + [\frac{(t^2 + 2tΔt +(Δt)^2}{1+t + Δt}-\frac{t^2}{1+t}] k} {Δt}"


"=\lim\limits_{Δt→0}\frac{[\frac{1+t-1-t-Δt} {(1+t) (1+t+ Δt)}] i + [\frac{(1+t)-t(1+t+Δt)}{(1+t)(1+t + Δt)} ]j + [\frac{(t+1)(t^2+2tΔt+(Δt)^2) -t^2(1+t+Δt)} {(1+t)(1+t + Δt)}] k} {Δt}"


"=\lim\limits_{Δt→0}\frac{[\frac{-Δt} {(1+t) (1+t+ Δt)}] i + [\frac{Δt} {(1+t)(1+t + Δt)} ]j + [\frac{ Δt (2tΔt+t^2+2t)} {(1+t)(1+t + Δt)}] k} {Δt}"


"=\lim\limits_{Δt→0}\frac{[Δt (\frac{-1} {(1+t) (1+t+ Δt)}) ] i + [ Δt(\frac{1} {(1+t)(1+t + Δt)}) ]j + [ Δt(\frac{ 2tΔt+t^2+2t} {(1+t)(1+t + Δt)}) ] k} {Δt}" Factorize Δt outside so that we have;


"= \lim\limits_{Δt→0}Δt\frac{ [ \frac{-1} {(1+t) (1+t+ Δt)} i + \frac{1} {(1+t)(1+t + Δt)} j + \frac{ 2tΔt+t^2+2t} {(1+t)(1+t + Δt)} k]} {Δt}" and it cancels out, remaining with;



"= \lim\limits_{Δt→0}{ \frac{-1} {(1+t) (1+t+ Δt)} i + \frac{1} {(1+t)(1+t + Δt)} j + \frac{ 2tΔt+t^2+2t} {(1+t)(1+t + Δt)} k}" Now substitute the value of Δt


"So, therefore, 𝐫'(t)= {\frac{-1} {(1+2t+t^2)} i + \frac{1} {(1+2t+t^2)} j + \frac{ t^2+2t} {(1+2t+t^2)} k}"